OpenStudy (anonymous):

Find the rate of change of q with respect to p when: p=20/(q^2+5) I get 40q'/(y^2+5)^2 but alas, my answer choices are -10/(qp^2), -5/(q^2p), -10/(qp), -5/(qp)^2. I think it's the last one but i'm not sure.

OpenStudy (anonymous):

why do you have a "y" in your answer?

OpenStudy (anonymous):

sorry it should say q

OpenStudy (anonymous):

those are your choices? wow.

OpenStudy (anonymous):

you are looking for \[q'\] right?

OpenStudy (anonymous):

yeah, I really hate this program we use it simplifies it in odd ways to trick you.

OpenStudy (anonymous):

lets try this and see if it works

OpenStudy (anonymous):

yes, with respect to p

OpenStudy (anonymous):

start with \[p=\frac{20}{q^2+5}\] \[p(q^2+5)=20\] \[pq^2+5p=20\] now take the derivative of q wrt p using implicit differentiation

OpenStudy (anonymous):

you get \[q^2+2pqq'=5=20\] and now solve for \[q'\]

OpenStudy (anonymous):

ok that is a typo let me write it correctly

OpenStudy (anonymous):

you get \[q^2+2pqq'+5=0\]

OpenStudy (anonymous):

so \[2pqq'=-5-q^2\] \[q'=\frac{-5-q^2}{2pq}\] is that one of the answers?

OpenStudy (anonymous):

no, I was thinking it was the last one. There was also a 'none of theese' answer but I selected it and got it wrong

OpenStudy (anonymous):

ok lets try this \[p=\frac{20}{q^2+5}\] \[\frac{20}{p}=q^2+5\] \[q^2=\frac{20}{p}+5\] and now take the derivative

OpenStudy (anonymous):

get \[2qq'=-\frac{20}{p^2}\] \[q'=-\frac{10}{qp^2}\] is that one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well maybe it is right. one can only hope

OpenStudy (anonymous):

It was right! Thank you. Can you help me with one more?